OpenStudy (anonymous):

show that eq has atleast one root in the given interval cos2x+sinx=0 , [-pi/2,pi/2]

OpenStudy (anonymous):

1-2sin^2x+sinx=0 2sin^2x-sinx-1=0 2sinx(sinx-1) +1(sinx-1)=0 sinx=1 or sinx=-1/2 x=pi/2 or x=-pi/6

OpenStudy (amistre64):

cos(2x) + sin(x) = 0 cos(2x) = sin(-x) cos^2 - sin^2 +sin(x) = 0 (1 - sin^2) +sin^2 + sin(x) = 0 1+sin(x) = 0 sin(x) = -1 x = -pi/2

OpenStudy (amistre64):

ack!!!....

OpenStudy (amistre64):

really need to do these on paper ;)

OpenStudy (anonymous):

\[cos(-\pi) = -1\] \[sin (-\pi/2) = -1\] \[cos(\pi) = -1\] \[sin(\pi/2) = 1\] So at least it will be 0 at \(\pi/2\). Another easy way would be if you found a positive and a negative then the intermediate value theorem would tell you you had a root somewhere in there.

OpenStudy (anonymous):

It doesn't specifically ask you to find it